Paired Exercise 14.2 Incorrect. Recall that there are exceptions to the solubility rules. Of the following substances, which ones are generally soluble in water? Cu(NO3)2 O AgNO3 OFe(OH)3 Ba3(PO4)2 SHOW HINT
media%2F215%2F215e579a-4b37-47ba-b287-02
Answer #1

14.2

Cu NO3)2 is soluble in water since almost all compounds containing NO_3^- ions are soluble as stated in solubility rules.

MgO is not soluble in water.

NH_4C_2H_3O_2 is soluble in water as it contains NH_4^+ ion.

AgNO_3 is soluble in water. Even though silver salts are not generally soluble, this is an exception to the rule.

Fe(OH)_3 is not soluble in water since most hydroxides of transition metal are not soluble.

Ba_3(PO_4)_2 is not soluble in water as most phosphates are not soluble according to solubility rules.

Hence, among the given option only NH_4C_2H_3O_2, AgNO_3 \ and \ Cu(NO_3)_2 are soluble in water.

14.14

(a)

The mass/volume percent is calculated using the following formula:

mass/volume \ \% = \frac{mass \ of \ solute \ in \ gm }{\ volume \ of \ solution \ in \ ml}\times 100 \%

Here, we have table sugar as solute, mass of table sugar = 161.2 gm

Voluem of the solution = 284.8 mL

hence,

mass/volume \ \% = \frac{161.2 \ g}{\ 284.8 ml}\times 100 \% = 56.6 \%

Hence, the mass by volume percent for the resulting solution is 56.6%.

(b)

Volume of solution = 75.5 ml

Mass of solute, methanol = 31.0 gm

mass/volume \ \% = \frac{31.0 \ g}{\ 75.5 ml}\times 100 \% = 41.06 \%

hence, the mass/volume percent for the resulting solution is 41.06%.